# RAMer: Reconstruction-based Adversarial Model for Multi-party Multi-modal Multi-label Emotion Recognition

### v1

@@ 5.1 Experimental setup
We evaluated RAMer on three multimodal, multi-label benchmark datasets: MEmoR Shen et al. (2020), a multi-party conversation dataset that includes personality, and CMU-MOSEI Zadeh et al. (2018) and $M^{3}$ Zhao et al. (2022), which are dyadic conversation datasets that do not include personality information. The evaluation is conducted under the protocols of these datasets. For CMU-MOSEI and $M^{3}$, we employed four commonly used evaluation metrics: Accuracy (Acc), Micro-F1, Precision (P), and Recall (R). Due to data imbalance in MEmoR, we followed the benchmark’s protocol and used Micro-F1, Macro-F1, and Weighted-F1 metrics.
Table 1: Performance comparison on the MEmoR dataset under primary and fine-grained settings.
With various modality combinations (visual($v$), acoustic($a$)), textual($t$), personality($p$)).
@@
Table 2: Performance Comparison on CMU-MOSEI dataset.
@@
Table 3: Performance Comparison on the M3ED dataset
@@
For the MEmoR dataset, we compare RAMer with multi-party conversation baselines, including MDL, MDAE Ngiam et al. (2011b), BiLSTM+TFN Zadeh et al. (2017), BiLSTM+LMF Liu et al. (2018b), DialogueGCN Ghosal et al. (2019), DialogueCRN Hu et al. (2021), and AMER Shen et al. (2020). We also evaluate its robustness against recent models designed for dyadic conversations, such as CARAT Peng et al. (2023) and TAILOR Zhang et al. (2022). For the CMU-MOSEI and $M^{3}$ datasets, we test three categories of methods. 1) Classic methods. CC Read et al. (2011), which concatenates all available modalities as input for binary classifiers. 2) Deep-based methods. ML-GCN Chen et al. (2019), using Graph Convolutional Networks to map label representations and capture label correlations. 3) Multi-modal multi-label methods. These include MulT Tsai et al. (2019) for cross-modal interactions, MISA Hazarika et al. (2020) for learning modality-invariant and modality-specific features, and methods like MMS2S Zhang et al. (2020), HHMPN Zhang et al. (2021a), TAILOR Zhang et al. (2022), AMP Ge et al. (2023), and CARAT Peng et al. (2023).
@@
Figure 3: (a) and (b), the t-SNE visualization of specific and common embedding without/with adversarial training. The red, green, and blue colors represent textual(t), visual(v), and acoustic(a) modalities respectively. Dark colors correspond to specific parts, and light colors denote common parts. (c) and (d), the t-SNE visualization of reconstruction embedding without/with RN and CLN. Different colors indicate different modalities and different saturation represents different emotions.
5.2 Comparison with the state-of-the-arts
@@
1) On the MEmoR dataset, RAMer outperforms all baselines by a significant margin. While TAILOR achieves a high weighted-F1 score in the fine-grained setting, its overall performance is weaker due to biases toward frequent and easier-to-recognize classes. RAMer consistently delivers strong results across all settings, demonstrating its ability to learn more effective representations. 2) On the CMU-MOSEI and M3ED datasets, RAMer surpasses state-of-the-art methods on all metrics except recall, which is less critical compared to accuracy and Micro-F1 in these contexts. 3) Deep-based methods outperform classical ones, highlighting the importance of capturing label correlations for improved classification performance. 4) Multimodal methods like HHMPN and AMP significantly outperform the unimodal ML-GCN, emphasizing the necessity of multimodal interactions. 5) Models optimized for dyadic conversations, such as CARAT, experience a notable performance drop in multi-party settings with incomplete modalities. In contrast, RAMer excels in both scenarios, achieving substantial improvements in Macro-F1 scores on the MEmoR dataset, outperforming CARAT by 0.178 and 0.209, respectively.
@@ 5.3 Ablation Study
  • The specificity and commonality enhance MMER performance. Variants (1), (2), and (3) exhibit an approximately 0.05 decrease in Micro-F1 performance compared to variant (11). This indicates that jointly learning specificity and commonalities yields superior performance, underscoring the importance of capturing both modality-specific specificity and shared commonality.
@@ 5.4 Qualitative Analysis
5.4.1 Visualization of Learned Modality Representations.
To evaluate the effectiveness of reconstruction-based adversarial training in generating distinguishable representations, we used t-SNE to visualize the commonality representations $\bm{C}_{\left\{v,a,t\right\}}$ and specificity representations $\bm{S}_{\left\{v,a,t\right\}}$ learned in the aligned CMU-MOSEI dataset. In figure 3(a), without adversarial training, specificity and commonality are loosely separated, but their distributions overlap in certain areas, such as the lower-right corner. In contrast, Figure 3(b) shows that with adversarial training, commonality and specificity are clearly separated in latent subspaces, forming distinct boundaries and effectively distinguishing emotions across modalities. Figure 3(c) demonstrates that without the reconstruction net (RN) and contrastive learning net (CLN), representations of different modalities are distinguishable, but emotion labels within the same modality remain intermixed. In contrast, Figure 3(d) shows that embeddings of both different modalities and labels are distinctly separable, highlighting that the reconstruction-based module effectively enhances modal specificity. Overall, RAMer accurately captures both the commonality and specificity of different modalities.
5.4.2 Visualization of Modality-to-Label Correlations.
@@ 5.4.3 Case Study
To demonstrate RAMer’s robustness in complex scenarios, Figure 5 shows an example of multi-party emotion recognition on MEmoR dataset where specific target persons have incomplete modality signals. The top three rows display different modalities from a video clip, segmented semantically with aligned multi-modal signals. Key observations include: 1) The target moment requires recognizing emotions for both the speaker (e.g., Howard) and non-speakers (e.g., Penny and Leonard). While the speaker typically has complete multi-modal signals, non-speakers often lack certain modalities. TAILOR, relying on incomplete modalities, produced partial predictions as its self-attention mechanisms struggled to align labels with missing features. 2) Limitations of single or incomplete modalities. A single modality, such as text, is often insufficient for accurate emotion inference (e.g., only Howard’s Joy is detectable from text alone). Although CARAT attempts to reconstruct missing information, it fails to capture cross-modal commonality, leading to incorrect predictions. 3)The importance of inter-person interactions and external knowledge is evident in emotion recognition, where inter-person attention enables individuals with incomplete modalities to gain supplementary information from others. Moreover, integrating external knowledge, such as personality traits, enhances emotion reasoning across participants and contexts, emphasizing the synergy between user profiling and emotion recognition. Experimental results validate that RAMer demonstrates superior robustness and effectiveness in these challenging, real-world scenarios.

### v2

@@ 5.1 Experimental Setup
We evaluated RAMer on three benchmark datasets: MEmoR Shen et al. (2020), a multi-party conversation dataset that includes personality, and CMU-MOSEI Zadeh et al. (2018) and $M^{3}$ED Zhao et al. (2022), which are dyadic conversation datasets that do not include personality information. The evaluation is conducted under the protocols of these datasets. For CMU-MOSEI and $M^{3}$ED, we employed four commonly used evaluation metrics: Accuracy (Acc), Micro-F1, Precision (P), and Recall (R). For MEmoR, we followed the benchmark’s protocol and used Micro-F1, Macro-F1, and Weighted-F1 metrics.
@@
Table 1: Performance comparison on the MEmoR dataset under primary and fine-grained settings. With various modality combinations (visual($v$), acoustic($a$), textual($t$), personality($p$)).
@@
Table 2: Performance Comparison on CMU-MOSEI dataset.
@@
For the MEmoR dataset, we compare RAMer with multi-party conversation baselines, including MDL, MDAE Ngiam et al. (2011b), BiLSTM+TFN Zadeh et al. (2017), BiLSTM+LMF Liu et al. (2018b), DialogueGCN Ghosal et al. (2019), DialogueCRN Hu et al. (2021), and AMER Shen et al. (2020). We further assess its robustness against recent dyadic models, including CARAT Peng et al. (2023) and TAILOR Zhang et al. (2022). For the CMU-MOSEI and $M^{3}$ED datasets, we test three categories of methods. 1) Classic methods. CC Read et al. (2011), which concatenates all available modalities as input for binary classifiers. 2) Deep-based methods. ML-GCN Chen et al. (2019), using Graph Convolutional Networks to map label representations and capture label correlations. 3) Multi-modal multi-label methods. These include MulT Tsai et al. (2019) for cross-modal interactions, MISA Hazarika et al. (2020) for learning modality-invariant and modality-specific features, and methods like MMS2S Zhang et al. (2020), HHMPN Zhang et al. (2021a), TAILOR Zhang et al. (2022), AMP Ge et al. (2023), and CARAT Peng et al. (2023).
@@
Figure 3: t-SNE visualizations of modality embeddings. (a)(b): Specificity and commonality features without/with adversarial training. Color indicates modality (textual, visual, acoustic); saturation differentiates specificity (dark) and commonality (light) components. (c)(d): Reconstruction embeddings without/with RN and CLN, where hue denotes modality and saturation encodes emotion.
5.2 Comparison with the State-of-the-Art
@@
1) On the MEmoR dataset, RAMer outperforms all baselines by a significant margin. While TAILOR achieves a high weighted-F1 score in the fine-grained setting, its overall performance is weaker due to biases toward frequent and easier-to-recognize classes. RAMer consistently delivers strong results across all settings, demonstrating its ability to learn more effective representations. 2) On the CMU-MOSEI and M3ED datasets, RAMer surpasses state-of-the-art methods on all metrics except recall, which is less critical compared to accuracy and Micro-F1 in these contexts. 3) Deep-based methods outperform classical ones, highlighting the importance of capturing label correlations for improved classification performance. 4) Multimodal methods like HHMPN and AMP significantly outperform the unimodal ML-GCN, emphasizing the necessity of multimodal interactions. 5) Models optimized for dyadic conversations, such as CARAT, experience a notable performance drop in multi-party settings with incomplete modalities. In contrast, RAMer excels in both scenarios, achieving substantial improvements in Micro-F1 and Macro-F1 scores on the MEmoR dataset.
@@ 5.3 Ablation Study
  • The specificity and commonality enhance MMER performance. Variants (1), (2), and (3) show lower Micro-F1 than variant (11). This indicates that jointly learning specificity and commonality yields superior performance, underscoring the importance of capturing both modality-specific specificity and shared commonality.
@@ 5.4 Qualitative Analysis
5.4.1 Visualization of Learned Modality Representations
To evaluate the effectiveness of reconstruction-based adversarial training, we use t-SNE to visualize the commonality and specificity representations learned in the aligned CMU-MOSEI dataset. In figure 3(a), without adversarial training, specificity and commonality are loosely separated, but their distributions overlap in certain areas, such as the lower-right corner. In contrast, Figure 3(b) shows clearer separation of commonality and specificity, forming distinct boundaries and effectively distinguishing emotions across modalities. Figure 3(c) demonstrates that without the reconstruction net (RN) and contrastive learning net (CLN), modality-wise embeddings are distinguishable, but emotion labels within the same modality remain intermixed. In contrast, Figure 3(d) shows clear separation across both modalities and emotions, indicating that the reconstruction-based module improves representation distinctiveness. Overall, RAMer accurately captures both the commonality and specificity of different modalities.
5.4.2 Visualization of Modality-to-Label Correlations
@@ 5.4.3 Case Study
To demonstrate RAMer’s robustness in complex scenarios, Figure 5 shows an example of MMER on the MEmoR dataset where specific target persons have incomplete modality signals. The top three rows display different modalities from a video clip, segmented semantically with aligned multi-modal signals. Key observations include: 1) The target moment requires recognizing emotions for both the speaker (e.g., Howard) and non-speakers (e.g., Penny and Leonard). While the speaker typically has complete multi-modal signals, non-speakers often lack certain modalities. TAILOR, limited by missing modalities, yields partial predictions as its self-attention mechanism struggles to align labels with missing features. 2) Limitations of single or incomplete modalities. A single modality, such as text, is often insufficient for accurate emotion inference (e.g., only Howard’s Joy is detectable from text alone). Although CARAT attempts to reconstruct missing information, it fails to capture cross-modal commonality, leading to incorrect predictions. 3)Inter-person interactions and external knowledge (e.g., personality traits) play important roles. Inter-person attention helps compensate for missing data, while personality-aware reasoning improves emotion inference across participants, highlighting the synergy between user profiling and emotion recognition. Experimental results demonstrate that RAMer achieves superior robustness and effectiveness in complex real-world scenarios.